(x-4y-9)dx+(4x+y-2)


1
Expert's answer
2021-09-09T23:59:29-0400

Let us solve the differential equation "(x-4y-9)dx+(4x+y-2)dy=0."


Firstly, consider the system "\begin{cases}x-4y-9=0 \\ 4x+y-2=0\end{cases}" which is equivalent to "\begin{cases}x-4y-9=0 \\ 16x+4y-8=0\end{cases}" and to "\begin{cases}x-4y-9=0 \\ 17x-17=0\end{cases}." It follows that "\begin{cases} x=1 \\ y=-2\end{cases}."


Let us use the transformation "x=u+1,\ y=v-2." Then we get "(u-4v)du+(4u+v)dv=0."

Also let "u=tv," and hence "du=tdv+vdt." It follows that "(tv-4v)(tdv+vdt)+(4tv+v)dv=0."

Since for "v=0" we have that "y=-2" is not a solution of the differential equation, let us divide both sides by "v:" "(t-4)(tdv+vdt)+(4t+1)dv=0." The last equation is equivalent to "(t^2-4t)dv+(t-4)vdt+(4t+1)dv=0," and hence to "(t^2+1)dv=-(t-4)vdt."

It follows that "-\frac{dv}{v}=\frac{(t-4)dt}{t^2+1}," and consequently "-\int\frac{dv}{v}=\int\frac{(t-4)dt}{t^2+1}." We get that "-\ln|v|=\frac{1}{2}\int\frac{2tdt}{t^2+1}-4\int\frac{dt}{t^2+1}=\frac{1}{2}\ln(t^2+1)-4\arctan t+\ln |C|," and hence

"8\arctan t=\ln(t^2+1)+2\ln|v|+\ln |C_1|=\ln|(t^2+1)v^2C_1|."

It follows that "(t^2+1)v^2C_1=e^{8\arctan t}."

Since "t=\frac{u}{v}=\frac{x-1}{y+2}" and "v=y+2," we have that

"((\frac{x-1}{y+2})^2+1)(y+2)^2C_1=e^{8\arctan \frac{x-1}{y+2}}."

We conclude that the general solution is the following:

"((x-1)^2+(y+2)^2)C_1=e^{8\arctan \frac{x-1}{y+2}}."